solve by substitution method
4x-3y=2
12x=6+9y
thanks very much for your help

Answer by xcentaur(357)

4x-3y=2......[1]
12x-9y=6
dividing by 3,
4x-3y=2......[2]


As you can see,equation one and two are both the same. for solving a pair of equations containing two variables, two seperate equations are required.
Please check your question and post again


-Prabhat